OpenStudy (anonymous):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1)

OpenStudy (aaronq):

q^2=1/2500 q=(1/2500)^(1/2) p=1-q 2pq= frequency of heterozygotes 2pq x 2500 = number of heterozygotes in every 2500 people so your answer will be "x in every 2500" You can't find the absolute answer unless you know the whole population